# Energy Window Augmented Plane Waves Approach to Density Functional Theory

### v5

@@ III.3 Expansion of basis set during convergence iteration
In many case it is worthwhile to use small basis sets at the beginning of the convergence loop and larger basis sets later (see [26] and Appendix C). In which case we need to add wavefunctions $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)$ with $\mathbf{G}_{min}^{n}<\mathbf{G}<\mathbf{G}_{max}^{n}$. These can be added to the main procedure in the same way as described in Section III.2.
@@ III.4 Partial basis (compatibility with $BEE-\left(LA\right)^{2}PW$ ideas [18])
III.5 EWAPW vs. Energy Window Linearized Plane Waves (EWLAPW) and multi-radius options [26]
The EWLAPW basis is described in Appendix A while multi radius options are described in Appendix B (see also ref. [26]). While these methods are slightly advantageous over EWAPW - in that they are slightly more accurate; the added complexity does not seem to be worth the minor improvement. To see this, we notice that the KS Hamiltonian, $-\nabla^{2}+V_{KS}^{n+1}\left(\mathbf{r}\right)$, has no singularities on the MT sphere radius. As such the eigenwavefunctions of the KS Hamiltonian should have no singularities near the surface of the MT sphere. Since the basis wavefunctions proposed in EWAPW are near eigenstates they should have no large singularities near the surface of the MT spheres and should be nearly continuously differentiable even though the constituent wavefunctions given in Eq. (10) have derivative discontinuities. As such adding $\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right)$ or varying the radius $S_{\alpha}$ to make the constituent wavefunctions continuously differentiable would not make much of a difference. From another perspective it is true that the scaling of the linearization error improves, we have that the linearization error changes:
@@
however it is already very small as the basis is nearly complete see Eq. (15) and the discussion below when the bands are well split into multiple windows so the improvement of multi radius options and EWLAPW options is minor.
@@ IV Energy Window Full Potential Augmented Plane Waves (EWFAPW)
Here we would like to do full potential calculations for the EWAPW basis, or EWFAPW. We break everything into simpler pieces - which are easier to code and reuse many pieces of existing codes - which should be evident by comparing this calculation with those of refs. [33, 29, 27, 30, 31, 28, 32, 34] for FLAPW. Because the wavefunction and its derivative are nearly continuous near the MT sphere (see Section III.5) it is reasonably appropriate to use the Hamiltonian from FLAPW (Eq. (16) without any boundary terms) appropriate to LAPW basis set, however the small corrections to the Hamiltonian due to discontinuity of derivatives of the wavefunction at the MT spheres have been included for accuracy in the form of a boundary term. This correction term comes about because the Laplacian term in the Hamiltonian is not self adjoint when the wavefunctions considered have derivative discontinuities. This may be corrected with a boundary term [22, 2]. Furthermore the small discontinuity of the wavefunction at the MT spheres (due to angular momentum truncation) can also be included [22] in the Hamiltonian. Parts that are not described namely the Weinert method [35], total energy as well as how to obtain $V_{KS}^{n}$ from $\rho^{n}\left(\mathbf{r}\right)$ go over verbatim in our construction from those corresponding to LAPW basis sets, that is that part of FLAPW code can be reused directly [33, 29, 27, 30, 31, 28, 32, 34].
@@ IV.2.2 MT Contribution
We now introduce the matrix of Hamiltonian radial overlaps:
@@ IV.2.3 Boundary Contribution
×$\times\left[\frac{4\pi}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}^{\prime}\right)\cdot\mathbf{r}_{\alpha}\right)Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}^{\prime}}\right)i^{l}\frac{\partial}{\partial r}j_{l}\left(\left|\mathbf{k}+\mathbf{G}^{\prime}\right|r\right)_{r=S_{\alpha}}-A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}\left(E_{i,n}^{V}\right)\frac{\partial}{\partial r}u_{l\alpha}^{n}\left(E_{i,n}^{V},r\right)_{r=S_{\alpha}}\right]$ (48)
@@ V Discussions & Conclusions
In this work we introduced a new method to solve the KS problem once the form of the correlation and exchange piece of the KS potential (be it within the Local Density Approximation (LDA), Generalized Gradient Approximation (GGA), metaGGA etc. [36, 39, 40, 37, 38]) has been fixed. In our method, EWAPW, we use the fact that solving the KS equations is an iterative process and use previous iteration’s eigenstate and eigenvalue information to produce a new basis for each iteration of the solution of the KS equations (see Section II.1). We have argued that no LO/lo or extensions (such as semicore LO, HDLO, HELO…) would significantly improve the accuracy of the solution as we have many linearization energies per occupied band (an a few for the unoccupied ones). As such we have created a competitor for LAPW+LO (as well as LAPW+HDLO, LAPW+HELO, LAPW+HDLO+HELO, etc.). However the basis set size we use is significantly smaller then LO type extensions as we have one or even less (see Section III.4) eigenstate per reciprocal lattice vector $\mathbf{G}$. This is an important advantage; indeed typical DFT calculations use 80-100 $\mathbf{G}$ vectors per atom [9, 18], while LO in the $s$, $p$, $d$, $f$ or even $g$ channels gives an additional 16-25 additional basis functions per LO band, so if we have LAPW+LO+HDLO+HELO($\times$3) [7, 18, 23] (where the ($\times$3) indicates three HELO bands are introduced) which is likely to have comparable accuracy to our methods, as it uses a comparable number of radial wavefunctions per atom, we can easily need to generate an additional 80-125 basis functions per atom. Doubling the basis size increase work load by a factor of ~eight so EWAPW is highly advantageous. Furthermore the Hamiltonian and overlap matrices for EWAPW are nearly diagonal and sparse giving additional speedup. In other words we have produced a basis with the efficiency of the APW+lo basis but with accuracy of LAPW+LO+HDLO+HELO.
@@
In this paper we focused on a specific method EWAPW, however as we discussed there are many variations of this method BEE-EWAPW, EWLAPW, multi radius options. Furthermore there are many slightly different ways of choosing the windowing function. While we have argued that there will not be a very significant change between all these methods, with respect to the accuracy of the solution of the KS problem, in the future it would still be of great interest to optimize the time requirements and the accuracy with respect to all these variations of the main method. Also in the future it would be of interest to combine these methods (EWAPW and related) with Hartree Fock as in DFT+U [41], or Dynamical Mean Field Theory, DFT+DMFT [41], or Gutzwiller Approximation, DFT+GA [42] to obtain more accurate results for strongly correlated systems. It would also be of interest to include magnetism, superconductivity and relativistic effects for enhanced accuracy. Besides expanding the methods it would also be of interest to expand applications, namely forces (including the Pulay contribution - which we believe to be very small in this case as the basis is nearly complete), pressure and stress [43]. As well as introduce DFT perturbation theory and study phonons with various methods [43]. This can be done for a large variety of options of EWAPW for DFT, DFT+U, DFT+DMFT, DFT+GA. We would also like to expand EWAPW to Energy Window Muffin Tin Orbitals (EWMTO [44]) and to energy window pseudopotential methods. This would greatly increase the applicability of energy window methods.
@@ Appendix A EWLAPW
$\Psi_{\mathbf{k}\mathbf{G}}^{Ln}\left(E\right)=\left\{\begin{array}[]{cc}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&\mathbf{r}\in IR\\ \left[a_{\mathbf{k}+\mathbf{G}}^{Ll\alpha n}\left(E\right)u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{Ln}\left(E\right)\right)+b_{\mathbf{k}+\mathbf{G}}^{Ll\alpha n}\left(E\right)\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{Ln}\left(E\right)\right)\right]Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (55)
@@ Appendix B Multi Radius options
Appendix C Adjusting the basis set size with iteration (review)

### v6

@@ III.3 Expansion of basis set during convergence iteration
In many case it is worthwhile to use small basis sets at the beginning of the convergence loop and larger basis sets later (see [26] and Appendix D). In which case we need to add wavefunctions $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)$ with $\mathbf{G}_{min}^{n}<\mathbf{G}<\mathbf{G}_{max}^{n}$. These can be added to the main procedure in the same way as described in Section III.2.
@@ III.4 Partial basis (compatibility with $BEE-\left(LA\right)^{2}PW$ ideas [18])
III.5 EWAPW vs. Energy Window Linearized Augmented Plane Waves (EWLAPW), multi-radius [26] and Continuous Interpolation Energy Window Augmented Plane Waves (CEWAPW).
The EWLAPW basis is described in Appendix A while multi radius options are described in Appendix B (see also ref. [26]) and CEWAPW is described in Appendix C. While these methods are slightly advantageous over EWAPW - in that they are slightly more accurate; the added complexity does not seem to be worth the minor improvement. To see this, we notice that the KS Hamiltonian, $-\nabla^{2}+V_{KS}^{n+1}\left(\mathbf{r}\right)$, has no singularities on the MT sphere radius. As such the eigenwavefunctions of the KS Hamiltonian should have no singularities near the surface of the MT sphere. Since the basis wavefunctions proposed in EWAPW are near eigenstates they should have no large singularities near the surface of the MT spheres and should be nearly continuously differentiable even though the constituent wavefunctions given in Eq. (11) have derivative discontinuities. As such adding $\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right)$ or varying the radius $S_{\alpha}$ to make the constituent wavefunctions continuously differentiable would not make much of a difference. From another perspective it is true that the scaling of the linearization error improves, we have that the linearization error changes:
@@
however it is already very small as the basis is nearly complete see Eq. (17) and the discussion below when the bands are well split into multiple windows so the improvement of multi radius, CEWAPW and EWLAPW options is minor. We note that continuous interpolation, multi radius and linearization options are combinable with each other.
@@ III.7 Generating the energy windows
III.8 Reducing the number of angular momentum channels
Since the EWAPW basis is a basis of near eigenstates of the KS Hamiltonian, the combinations $\sum_{\mathbf{G}}\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)$ are given by:
$\sum_{\mathbf{G}}\frac{\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)=4\pi\sum_{l,m}i^{l}\left[\sum_{\mathbf{G}}\frac{\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}}{\sqrt{\Omega}}Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}}\right)j_{l}\left(\left|\mathbf{k}+\mathbf{G}\right|\left|\mathbf{r}-\mathbf{r}_{\alpha}\right|\right)\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}_{\alpha}\right)\right]Y_{lm}\left(\widehat{\mathbf{r-\mathbf{r_{\alpha}}}}\right)$ (29)
Whereby we have that
$\sum_{\mathbf{G}}\frac{\bar{z}_{\mathbf{k}\mathbf{G}}^{\nu n}}{\sqrt{\Omega}}Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}}\right)j_{l}\left(\left|\mathbf{k}+\mathbf{G}\right|\left|\mathbf{r}-\mathbf{r}_{\alpha}\right|\right)\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}_{\alpha}\right)\sim\delta_{l,l_{0}}+corrections$ (30)
Where $l_{0}$ is the dominant angular momentum channel of the band in question. From this we see that only a small number of angular momentum channels participate much like in lo/LO constructions so the truncation angular momentum $l_{max}$ may be reduced.
III.9 Efficiency of method
We have that the basis size for the EWAPW method is very comparable to that of APW but the construction of the KS Hamiltonian and charge density requires a large number of radial integrals see Eqs. (37) and (50) below. This limits the number of possible windows. However there are many mitigating factors: the angular momentum channels of the EWAPW wavefunctions can be kept small much like those of lo/LO wavefunctions furthermore for a specific $\mathbf{k}$ point one never needs to consider as many radial wavefunctions as the total number of bands that are considered relevant to the windowing construction (see Section III.7 for an example of relevant bands where the occupied and lowest $\mathfrak{M}$ unoccupied bands are relevant so not relevant to the window $\left[E_{N-1,n+1}^{U},E_{N,n+1}^{U}=+\infty\right)$). Furthermore for large scale calculations the diagonalization time is the dominant contribution to the computer time and should be very efficient with the EWAPW basis.
@@ IV Energy Window Full Potential Augmented Plane Waves (EWFAPW)
Here we would like to do full potential calculations for the EWAPW basis, or EWFAPW. We break everything into simpler pieces - which are easier to code and reuse many pieces of existing codes - which should be evident by comparing this calculation with those of refs. [33, 29, 27, 30, 31, 28, 32, 34] for FLAPW. Because the wavefunction and its derivative are nearly continuous near the MT sphere (see Section III.5) it is reasonably appropriate to use the Hamiltonian from FLAPW (Eq. (18) without any boundary terms) appropriate to LAPW basis set, however the small corrections to the Hamiltonian due to discontinuity of derivatives of the wavefunction at the MT spheres have been included for accuracy in the form of a boundary term. This correction term comes about because the Laplacian term in the Hamiltonian is not self adjoint when the wavefunctions considered have derivative discontinuities. This may be corrected with a boundary term [22, 2, 49, 50]. Furthermore the small discontinuity of the wavefunction at the MT spheres (due to angular momentum truncation) can also be included [22] in the Hamiltonian. Parts that are not described namely the Weinert method [35], total energy as well as how to obtain $V_{KS}^{n}$ from $\rho^{n}\left(\mathbf{r}\right)$ go over verbatim in our construction from those corresponding to LAPW basis sets, that is that part of FLAPW code can be reused directly [33, 29, 27, 30, 31, 28, 32, 34].
@@ IV.2.2 MT Contribution
We first note that
$=\int_{MT_{\alpha}}d\mathbf{r}\Psi_{\mathbf{k}j}^{n}{}^{*}\left(\mathbf{r}\right)E_{i,n}^{V}\Psi_{\mathbf{k}k}^{n}\left(\mathbf{r}\right)$ (47)
for some $i$. We now introduce the matrix of Hamiltonian radial overlaps:
@@ IV.2.3 Boundary Contribution
×$\times\left[\frac{4\pi}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}^{\prime}\right)\cdot\mathbf{r}_{\alpha}\right)Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}^{\prime}}\right)i^{l}\frac{\partial}{\partial r}j_{l}\left(\left|\mathbf{k}+\mathbf{G}^{\prime}\right|r\right)_{r=S_{\alpha}}-\bar{A}_{\mathbf{k}\mathbf{G}^{\prime}}^{lm\alpha n}\left(E_{i,n}^{V}\right)\frac{\partial}{\partial r}u_{l\alpha}^{n}\left(E_{i,n}^{V},r\right)_{r=S_{\alpha}}\right]$ (53)
@@ V Discussions & Conclusions
In this work we introduced a new method to solve the KS problem once the form of the correlation and exchange piece of the KS potential (be it within the Local Density Approximation (LDA), Generalized Gradient Approximation (GGA), metaGGA etc. [36, 39, 40, 37, 38]) has been fixed. In our method, EWAPW, we use the fact that solving the KS equations is an iterative process and use previous iteration’s eigenstate and eigenvalue information to produce a new basis for each iteration of the solution of the KS equations (see Section II.1). We have argued that no LO/lo or extensions (such as semicore LO, HDLO, HELO…) would significantly improve the accuracy of the solution as we have many linearization energies per occupied band (an a few for the unoccupied ones). As such we have created a competitor for LAPW+LO (as well as LAPW+HDLO, LAPW+HELO, LAPW+HDLO+HELO, etc.). However the basis set size we use is significantly smaller then LO type extensions as we have one or even less (see Section III.4) eigenstate per reciprocal lattice vector $\mathbf{G}$. This is an important advantage; indeed typical DFT calculations use 80-100 $\mathbf{G}$ vectors per atom for large MT spheres [9, 18], while LO in the $s$, $p$, $d$, $f$ or even $g$ channels gives an additional 16-25 additional basis functions per LO band, so if we have LAPW+LO+HDLO+HELO($\times$3) [7, 18, 23] (where the ($\times$3) indicates three HELO bands are introduced) which is likely to have comparable accuracy to our methods, as it uses a comparable number of radial wavefunctions per atom, we can easily need to generate an additional 80-125 basis functions per atom. Doubling the basis size increase work load by a factor of ~eight so EWAPW is highly advantageous. Furthermore the Hamiltonian and overlap matrices for EWAPW are nearly diagonal and sparse giving additional speedup. In other words we have produced a basis with the efficiency of the APW basis but with accuracy of LAPW+LO+HDLO+HELO. We would like to note that the efficiency of a DFT method is mostly determined by the efficiency of the diagonalization process and the efficiency of setting up the KS Hamiltonian, overlap matrix and charge distribution. While the EWAPW basis has the efficiency of diagonalization of APW or better (same basis size, but a sparser matrix) EWAPW takes comparable time to LAPW+LO+HDLO+HELO to set up the KS Hamiltonian and overlap matrices as well as the electron density.
@@
In this paper we focused on a specific method EWAPW, however as we discussed there are many variations of this method BEE-EWAPW, EWLAPW, multi radius CEWAPW options. Furthermore there are many slightly different ways of choosing the windowing function. While we have argued that there will not be a very significant change between all these methods, with respect to the accuracy of the solution of the KS problem, in the future it would still be of great interest to optimize the time requirements and the accuracy with respect to all these variations of the main method. Also in the future it would be of interest to combine these methods (EWAPW and related) with Hartree Fock as in DFT+U [41], or Dynamical Mean Field Theory [52], DFT+DMFT [41], or Gutzwiller Approximation, DFT+GA [42] to obtain more accurate results for strongly correlated systems. It would also be of interest to include magnetism, superconductivity and relativistic effects for enhanced accuracy. Besides expanding the methods it would also be of interest to expand applications, namely forces (including the Pulay contribution - which we believe to be very small in this case as the basis is nearly complete [48]), pressure and stress [43, 51]. As well as introduce DFT perturbation theory and study phonons with various methods [43]. This can be done for a large variety of options of EWAPW for DFT, DFT+U, DFT+DMFT, DFT+GA. We would also like to expand EWAPW to Energy Window Muffin Tin Orbitals (EWMTO [44]) and to energy window pseudopotential methods. This would greatly increase the applicability of energy window methods.
@@ Appendix A EWLAPW
$\Psi_{\mathbf{k}\mathbf{G}}^{Ln}\left(E\right)=\left\{\begin{array}[]{cc}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&\mathbf{r}\in IR\\ \left[a_{\mathbf{k}+\mathbf{G}}^{Ll\alpha n}\left(\mathcal{E}_{w}\left(E\right)\right)u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{Ln}\left(E\right)\right)+b_{\mathbf{k}+\mathbf{G}}^{Ll\alpha n}\left(\mathcal{E}_{w}\left(E\right)\right)\dot{u}_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{Ln}\left(E\right)\right)\right]Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (60)
@@ Appendix B Multi Radius options
Appendix C CEWAPW
CEWAPW is based on the following wavefunctions:
$\tilde{\Phi}_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)$ $=\sum_{i=3}^{N-1}\left[\frac{\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{i-1,n}^{CU}\right)\left(E_{i,n}^{CU}-E\right)+\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{i,n}^{CU}\right)\left(E-E_{i-1,n}^{CU}\right)}{E_{i,n}^{CU}-E_{i-1,n}^{CU}}\right]\cdot\left(\Theta\left(E-E_{i-1,n}^{CU}\right)-\Theta\left(E-E_{i,n}^{CU}\right)\right)$
$+\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{2,n}^{CU}\right)\Theta\left(E_{2,n}^{CU}-E\right)+\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{N-1,n}^{CU}\right)\Theta\left(E-E_{N-1,n}^{CU}\right)$ (66)
Where $E_{1,n}^{CU}=-\infty$, $E_{N,n}^{CU}=+\infty$ and $E_{i-1,n}^{CU}<E_{i,n}^{CU}$ $\forall i$. These wavefunctions continuously interpolate between the various wavefunctions $\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{i,n}^{CU}\right)$. One then replaces $\tilde{\Phi}_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)\leftrightarrow\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)$ in Section II.1 and proceeds to build a basis of near eigenstates. Alternatively one could use even more sophisticated interpolation schemes and write:
$\hat{\Phi}_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=\sum_{i=2}^{N-1}\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E_{i,n}^{CU}\right)\prod_{j\neq i}\frac{\left(E-E_{j,n}^{CU}\right)}{\left(E_{i,n}^{CU}-E_{j,n}^{CU}\right)}$ (67)
One then replaces $\hat{\Phi}_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)\leftrightarrow\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)$ in Section II.1 and proceeds to build a basis of near eigenstates. In general any linear interpolation method would do.
Appendix D Adjusting the basis set size with iteration (review)